Dec 4 (UT)

Observers: Wen-fai Fong

PI programs: Fong

00:00 UT -- Dome opened.

01:00 UT -- Focusing, stars have a large coma. Global offsets were causing comas. Lots of trouble-shooting. Found that by clearing global offsets it became fine. Also some pointing guide star camera issues that delayed by another 30 minutes. Lost ~2.5 targets.

03:52 UT -- Start 1st science target, airmass~1.1, seeing~0.7".

04:33 UT -- Acquire 2nd target.

04:41 UT -- Start 2nd science target, airmass~1.1, seeing~0.66".

06:30 UT -- Start 3rd science target, airmass 1.5, seeing~0.95"

07:16 UT -- Start 4th science target, airmass 1.2, seeing~0.99".

09:05 UT -- Start 5th science target, airmass 1.3, seeing~1.02".

09:30-09:35 UT -- Paused, lost guiding. Resumed after 5 min in partial clouds.

09:55 UT - Stopped script, lost guiding.

10:02 UT -- Resumed script, relatively poor conditions.

10:00-10:30 UT -- In and out of clouds, paused script several times and eventually gave up.

10:30 UT -- Recollimating telescope, in thicker clouds.

10:40 UT -- Standard star G191-B2B (in thick'ish clouds), airmass~1.3, 0.8" seeing.

11:00 UT -- Stopped for weather after some trouble guiding. Thick clouds.

11:27 UT -- Start 6th science target, airmass ~1.1, 0.9" seeing. Clearing skies.

12:27 UT -- Start final science target, airmass~1.1, 0.79" seeing. Clear skies.

13:14 UT -- End of observing.

-- %USERSIG{EiichiEgami - 2015-12-02}%

Comments


 
Topic revision: r6 - 04 Dec 2015, EiichiEgami
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback